# Nuisance-free Automatic Ground Collision Avoidance System Design: Merging Exponential-CBF and Adaptive Sliding Manifolds

### v1

@@ III-A Nuisance-free Intervention Design
$\underbrace{\mathcal{L}^{2}_{f}b(h)+\mathcal{L}_{g}\mathcal{L}_{f}b(h)q_{\text{GCAS}}}_{\begin{subarray}{c}\ddot{h}\end{subarray}}+\underbrace{\kappa\eta_{b}(h)}_{\begin{subarray}{c}k_{2}\dot{h}+k_{1}h\end{subarray}}\geq 0$ (31)
The system also represents the second-order altitude dynamics, and the state-space form of the obtained system is given in Eq. (32), with a new state vector defined as $\bm{\xi}=[\xi_{1}\hskip 4.26773pt\xi_{2}]^{T}$.
$\begin{bmatrix}\dot{\xi_{1}}\\ \dot{\xi_{2}}\\ \end{bmatrix}=\underbrace{\begin{bmatrix}0&1\\ -k_{1}&-k_{2}\end{bmatrix}}_{\begin{subarray}{c}\Phi(\kappa)\end{subarray}}\begin{bmatrix}\xi_{1}\\ \xi_{2}\\ \end{bmatrix}$ (32)
@@
Then, the proposition in III.1 must hold for the equilibrium point to be an attractive node.
Proposition III.1.
@@
Lastly, the proposition in III.2 must hold for the system dynamics to have a damping ratio $\zeta=1$.
Proposition III.2.
@@
$\mathcal{S}=\{(\kappa(\bm{x}),\bm{\xi})|\sigma(\kappa(\bm{x}),\bm{\xi})=0\}$ (33)
@@
$\sigma(\kappa(\bm{x}),\bm{\xi})=\xi_{2}+\mu\xi_{1}$ (34)
where $\bm{x}=[\phi,\theta,V_{T}]$ and $\mu=k_{1}/k_{2}$ is the adaptive design parameter. The given definition of the adaptive sliding manifold implies that the sliding manifolds are principally functions of the gain $\kappa$, and the gain $\kappa$ is a function of the governing flight states, i.e. $\phi$, $\theta$, and $V_{T}$. Consequently, the adaptivity of the gain $\kappa$ with respect to $\bm{x}=[\phi,\theta,V_{T}]$ and $\mu=k_{1}/k_{2}$ directly leads to the adaptivity of sliding manifolds. For this purpose, the adaptive design of gain $\kappa$ is performed through an optimization framework, which is presented in the proceeding section.
III-A1 Optimization Framework
An optimization problem is set up to calculate the gain pair $k_{1}$ and $k_{2}$ in accordance with the timely and aggressive intervention requirements. To obtain the reference command shape, represented in Fig. 1, the principal objectives are determined as follows: (1) the integral of the pitch rate command within its applied time interval, denoted as $[t_{0},\hskip 4.26773ptt_{f}]$, (2) the maximum amplitude of the pitch rate command within its applied time interval, denoted as $[t_{0},\hskip 4.26773ptt_{f}]$, and (3) the norm of the difference between the minimum altitude of the aircraft and the DTED with buffer height, i.e., $h_{\text{DTED}}+h_{\text{buff}}$. Thereby, the optimization problem can be described as a multi-objective optimization problem, where the compact objective function is given by Eq. (35).
[…]
$\cdots+W_{3}\underbrace{|\min\big{(}h(t)\big{)}-(h_{\text{buff}}+h_{\text{DTED}})|}_{\begin{subarray}{c}J_{3}\end{subarray}}$
@@ III-A1 Optimization Framework
Remark III.1.
If Proposition III.1 and III.2 are substituted, the optimization problem is reduced to a single variable multi-objective optimization, where $k_{1}=k_{2}^{2}/4$ and $k_{1}\in[4,\infty)$.
@@
$\min_{k_{1},k_{2}\in\mathbb{R}_{>0}}$ $\sum^{3}_{i=1}W_{i}\hat{J}_{i}$ (36)
@@
$k_{1}\geq 4$
where $\hat{J}$ represent the normalized objective function between 00 and $1$.
Subsequently, the optimization is performed under various conditions by meshing $\phi$, $\theta$, and $V_{T}$ within the intervals $[-150^{\circ},\hskip 4.26773pt150^{\circ}]$ with an increment of $\Delta\phi=50^{\circ}$, $[-60^{\circ},\hskip 4.26773pt-10^{\circ}]$ with an increment of $\Delta\theta=10^{\circ}$, and $[200,\hskip 4.26773pt350]$ with an increment of $\Delta V_{T}=30$m/s, respectively. The other flight states are randomly assigned to ensure the aircraft dives. Finally, the optimization is carried out at each design point, assuming the ground is flat and at zero level, i.e., $h_{\text{DTED}}=0$m while keeping $h_{\text{buff}}=100$ m, and the calculated gains $k_{1}$ and $k_{2}$ are scheduled as functions of $\phi$, $\theta$, and $V_{T}$.
IV Flight Envelope Protection System Design
The flight envelope protection serves as an additional safety layer to ensure “do no harm” principle. For this purpose, the protection algorithms for the angle of attack, load factor, and bank angle are designed using CBF.
IV-A Angle of Attack/Load Factor Protection
@@
$b(\alpha)=\alpha_{\text{limit}}-\alpha$ (37)
where $\alpha_{\text{limit}}\in\mathbb{R}$. It is obvious that $b(\alpha)>0,\forall\alpha\in\mathbb{R}_{<\alpha_{\text{limit}}}$, and $b(\alpha)=0,\alpha=\alpha_{\text{limit}}$. Thus, the time derivative of the barrier function is $\dot{b}(\alpha)=-\dot{\alpha}$. For the sake of convenience, the load factor limit can be described in the form of an angle of attack limit, as presented in Eq. (38).
$\alpha^{n_{z}}_{\text{limit}}=\dfrac{n_{z_{\text{limit}}}mg}{\bar{q}_{\infty}SC_{z_{\alpha}}}$ (38)
where $C_{z_{\alpha}}$ denotes the z-axis force coefficient derivative with respect to the angle of attack. The most restrictive limit can then be selected between $\alpha_{\text{limit}}$ and $\alpha^{n_{z}}_{\text{limit}}$. The primary limit for the angle of attack is considered to be the stall angle of attack, $\alpha_{\text{stall}}$; therefore, the equivalent angle of attack limit can be expressed as given in Eq. (39).
$\alpha_{\text{limit}}=\min(\alpha_{\text{stall}},\alpha^{n_{z}}_{\text{limit}})$ (39)
@@
$\underbrace{-f(\alpha)-g(\alpha)q}_{\begin{subarray}{c}\mathcal{L}_{f}b(\alpha)+\mathcal{L}_{g}b(\alpha)q_{\text{cmd}}\end{subarray}}+\gamma_{\alpha}b(\alpha)\geq 0$ (40)
@@
$\min_{q_{\text{cmd}}\in\mathbb{R}}$ $\dfrac{1}{2}\big{(}q_{\text{cmd}}-q_{\text{GCAS}}\big{)}^{2}$ (41)
@@
$q_{\text{min}}\leq q_{\text{cmd}}\leq q_{\text{max}}$
@@
IV-B Bank Angle Protection
@@
$b(\phi)=\phi_{\text{limit}}-|\phi|$ (42)
where $\phi_{\text{limit}}\in\mathbb{R}$. It is obvious that $b(\phi)>0,\forall\phi\in\mathbb{R}_{<\phi_{\text{limit}}}$, and $b(\phi)=0,\phi=\phi_{\text{limit}}$. Thus, the time derivative of the barrier function is $\dot{b}(\phi)=-\text{sgn}(\phi)\dot{\phi}$. Additionally, since the main function of the bank angle protection is to conduct a bank-to-level maneuver, the limit of the bank angle, $\phi_{\text{limit}}$, is simply set to $0^{\circ}$.
@@
$\underbrace{-\text{sgn}(\phi)[f(\phi)+g(\phi)p]}_{\begin{subarray}{c}\mathcal{L}_{f}b(\phi)+\mathcal{L}_{g}b(\phi)p_{\text{GCAS}}\end{subarray}}+\gamma_{\phi}b(\phi)\geq 0$ (44)
where $\gamma_{\phi}\in\mathbb{R}_{>0}$ is the design parameter to be chosen properly. As a consequence, the final form of the bank angle protection formulation for generating a roll rate command is presented in Eq. (45).
$\min_{p_{\text{GCAS}}\in\mathbb{R}}$ $\dfrac{1}{2}\big{(}p_{\text{GCAS}}-p_{\text{pilot}}\big{)}^{2}$ (45)
s.t. $-\text{sgn}(\phi)[f(\phi)+g(\phi)p_{\text{GCAS}}]+\gamma_{\phi}b(\phi)\geq 0$
$p_{\text{min}}\leq p_{\text{GCAS}}\leq p_{\text{max}}$
The constructed formulation enables the Auto-GCAS to generate a roll rate command ($p_{\text{GCAS}}$) that closely follows the roll rate command of pilot ($p_{\text{pilot}}$) while adhering to the constraint for ground collision avoidance, i.e. $\phi=0^{\circ}$. Additionally, the generated roll rate command, $p_{\text{GCAS}}$, must remain within the interval $[p_{\text{min}}\hskip 4.26773ptp_{\text{max}}]$, considering the admissible and allowable roll rate limits depending on the aircraft. Since there does not exist an additional layer for the bank angle protection, the command $p_{\text{GCAS}}$ is equivalent to the command $p_{\text{cmd}}$.
Remark IV.1.
It is noteworthy that the bank angle must be within the interval, $-\pi\leq\phi\leq\pi$, by definition.
@@
Figure 4: General framework of the proposed method: (1) Safety layer including GCAS and FEP structures, (2) CAS including NDI, (3) Control allocation including INCA, and (4) A/C dynamics. The pilot commands $p_{\text{pilot}}$, $q_{\text{pilot}}$, and $r_{\text{pilot}}$, with roll and pitch rate commands subject to GCAS constraints. The pitch rate command is adjusted by the ECBF constraint, and $q_{\text{GCAS}}$ is generated if needed. This command is also limited by angle of attack and load factor protection, generating $q_{\text{cmd}}$ when required. Bank angle protection initiates a bank-to-level maneuver, generating $p_{\text{GCAS}}$ if $q_{\text{pilot}}-q_{\text{GCAS}}\neq 0$, indicating a recovery maneuver initiated; otherwise, the pilot’s roll rate command, $p_{\text{pilot}}$, is passed through.
V Results
The proposed Auto-GCAS design has been evaluated through three different assessments, (1) terrain collision avoidance scenarios, (2) an authority-sharing scenario, and (3) Monte Carlo simulations. Note that, for all the assessments, the buffer height is set to 100 meters.
V-A Terrain Collision Avoidance Assessments
Two distinct terrain collision avoidance scenarios are conceptualized: one at low altitude with randomized state variables and another at high altitude with randomized states variables. The term “randomized state variables” encompasses aerodynamic angles, Euler angles, angular rates, and true velocity, with altitude being explicitly excluded. A default terrain model, synthesized using MATLAB®’s peaks function, serves as the basis for scenario generation. Furthermore, it is assumed that the pilot is incapacitated (e.g. experiencing blackout), resulting in the absence of control inputs, such that angular rate commands are maintained at $0^{\circ}/\text{s}$. Finally, it is important to note that a terrain scan pattern should ideally be employed, given that the terrain is not flat but mountainous. However, the design of such a pattern falls beyond the scope of this study. Consequently, a default rectangular scan pattern is employed, extending 750 meters in height and 150 meters to each side (left and right) in width from the nose of the aircraft.
V-A1 Scenario-1: Low Altitude

### v2

@@ 3.1 Nuisance-free Intervention Design
$\underbrace{\mathcal{L}^{2}_{f}b(h)+\mathcal{L}_{g}\mathcal{L}_{f}b(h)q_{\text{GCAS}}}_{\begin{subarray}{c}\ddot{h}\end{subarray}}+\underbrace{\kappa\eta_{b}(h)}_{\begin{subarray}{c}k_{2}\dot{h}+k_{1}h\end{subarray}}\geq 0.$ (32)
The system also represents the second-order altitude dynamics, and the state-space form of the obtained system is given in Eq. (33), with a new state vector defined as $\bm{\xi}=[\xi_{1}\hskip 4.26773pt\xi_{2}]^{\mathrm{T}}$.
$\begin{bmatrix}\dot{\xi_{1}}\\ \dot{\xi_{2}}\\ \end{bmatrix}=\underbrace{\begin{bmatrix}0&1\\ -k_{1}&-k_{2}\end{bmatrix}}_{\begin{subarray}{c}\Phi(\kappa)\end{subarray}}\begin{bmatrix}\xi_{1}\\ \xi_{2}\\ \end{bmatrix}$, (33)
@@
Then, the proposition in 3.1 must hold for the equilibrium point to be an attractive node.
Proposition 3.1.
@@
Lastly, the proposition in 3.2 must hold for the system dynamics to have a damping ratio $\zeta=1$.
Proposition 3.2.
@@
$\mathcal{S}=\{(\kappa(\bm{x}),\bm{\xi})|\sigma(\kappa(\bm{x}),\bm{\xi})=0\},$ (34)
@@
$\sigma(\kappa(\bm{x}),\bm{\xi})=\xi_{2}+\mu\xi_{1},$ (35)
where $\bm{x}=[\phi\hskip 4.26773pt\theta\hskip 4.26773ptV_{T}]^{\mathrm{T}}$ and $\mu=k_{1}/k_{2}$ is the adaptive design parameter. The given definition of the adaptive sliding manifold implies that the sliding manifolds are principally functions of the gain $\kappa$, and the gain $\kappa$ is a function of the governing flight states, i.e. $\phi$, $\theta$, and $V_{T}$. Consequently, the adaptivity of the gain $\kappa$ with respect to $\bm{x}=[\phi\hskip 4.26773pt\theta\hskip 4.26773ptV_{T}]^{\mathrm{T}}$ and $\mu=k_{1}/k_{2}$ directly leads to the adaptivity of sliding manifolds. For this purpose, the adaptive design of gain $\kappa$ is performed through an optimization framework, which is presented in the proceeding section.
3.1.1 Optimization Framework
An optimization problem is set up to calculate the gain pair $k_{1}$ and $k_{2}$ in accordance with the timely and aggressive intervention requirements. To obtain the reference command shape, represented in Fig. 1, the principal objectives are determined as follows: (1) the integral of the pitch rate command within its applied time interval, denoted as $t_{0}\hskip 4.26773ptt_{f}]$, (2) the maximum amplitude of the pitch rate command within its applied time interval, denoted as $t_{0}\hskip 4.26773ptt_{f}]$, and (3) the norm of the difference between the minimum altitude of the aircraft and the DTED with buffer height, i.e., $h_{\text{DTED}}+h_{\text{buff}}$. Thereby, the optimization problem can be described as a multi-objective optimization problem, where the compact objective function is given by Eq. (36).
$J_{T}$ $=W_{1}\underbrace{\Big{[}-\int^{t_{f}}_{t_{0}}q_{\text{GCAS}}(t)dt\Big{]}}_{\begin{subarray}{c}J_{1}\end{subarray}}+W_{2}\underbrace{\big{[}-\max(q_{\text{GCAS}}(t))\big{]}}_{\begin{subarray}{c}J_{2}\end{subarray}}+W_{3}\underbrace{|\min\big{(}h(t)\big{)}-(h_{\text{buff}}+h_{\text{DTED}})|}_{\begin{subarray}{c}J_{3}\end{subarray}},$ (36)
@@ 3.1.1 Optimization Framework
Remark 3.1.
If Proposition 3.1 and 3.2 are substituted, the optimization problem is reduced to a single variable multi-objective optimization, where $k_{1}=k_{2}^{2}/4$ and $k_{1}\in[4,\infty)$.
@@
$k^{\star}_{1},k^{\star}_{2}=\operatorname*{argmin}_{k_{1},k_{2}\in\mathbb{R}_{>0}}$ $\sum^{3}_{i=1}W_{i}\hat{J}_{i}$ (37)
@@
$k_{1}\geq 4.$
where $\hat{J}$ represent the normalized objective function between 0 and $1$.
4 Flight Envelope Protection System Design
The flight envelope protection serves as an additional safety layer to ensure “do no harm" principle. For this purpose, the protection algorithms for the angle of attack, load factor, and bank angle are designed using CBF.
4.1 Angle of Attack/Load Factor Protection
@@
$b(\alpha)=\alpha_{\text{limit}}-\alpha,$ (38)
where $\alpha_{\text{limit}}\in\mathbb{R}$. It is obvious that $b(\alpha)>0,\forall\alpha\in\mathbb{R}_{<\alpha_{\text{limit}}}$, and $b(\alpha)=0\iff\alpha=\alpha_{\text{limit}}$. Thus, the time derivative of the barrier function is $\dot{b}(\alpha)=-\dot{\alpha}$. For convenience, the load factor limit can be expressed as an angle of attack limit (due to the bank-to-level requirement, i.e., $\phi=0^{\circ}$), as presented in Eq. (39).
$\alpha^{n_{z}}_{\text{limit}}=\dfrac{n_{z_{\text{limit}}}mg}{\bar{q}_{\infty}SC_{z_{\alpha}}},$ (39)
where $C_{z_{\alpha}}$ denotes the z-axis force coefficient derivative with respect to the angle of attack. To ensure safety, the equivalent angle of attack constraint is selected as the more restrictive value between the nominal stall angle $\alpha_{\text{limit}}$ and the load factor-based limit $\alpha^{n_{z}}_{\text{limit}}$. The primary limit for the angle of attack is considered to be the stall angle of attack, $\alpha_{\text{stall}}$; therefore, the equivalent angle of attack limit can be expressed as given in Eq. (40).
$\alpha_{\text{limit}}=\min(\alpha_{\text{stall}},\alpha^{n_{z}}_{\text{limit}}).$ (40)
@@
$\underbrace{-f(\alpha)-g(\alpha)q}_{\begin{subarray}{c}\mathcal{L}_{f}b(\alpha)+\mathcal{L}_{g}b(\alpha)q_{\text{cmd}}\end{subarray}}+\gamma_{\alpha}b(\alpha)\geq 0,$ (41)
@@
$u^{\star}=\operatorname*{argmin}_{q_{\text{cmd}}\in\mathbb{R}}$ $\dfrac{1}{2}\big{(}q_{\text{cmd}}-q_{\text{GCAS}}\big{)}^{2}$ (42)
@@
$q_{\text{min}}\leq q_{\text{cmd}}\leq q_{\text{max}}.$
@@
4.2 Bank Angle Protection
@@
$b(\phi)=\phi_{\text{limit}}-|\phi|,$ (43)
where $\phi_{\text{limit}}\in\mathbb{R}$. It is obvious that $b(\phi)>0,\forall\phi\in\mathbb{R}_{<\phi_{\text{limit}}}$, and $b(\phi)=0\iff\phi=\phi_{\text{limit}}$. Thus, the time derivative of the barrier function is $\dot{b}(\phi)=-\text{sgn}(\phi)\dot{\phi}$. Additionally, since the main function of the bank angle protection is to conduct a bank-to-level maneuver, the limit of the bank angle, $\phi_{\text{limit}}$, is simply set to $0^{\circ}$.
@@
$\underbrace{-\text{sgn}(\phi)[f(\phi)+g(\phi)p]}_{\begin{subarray}{c}\mathcal{L}_{f}b(\phi)+\mathcal{L}_{g}b(\phi)p_{\text{GCAS}}\end{subarray}}+\gamma_{\phi}b(\phi)\geq 0,$ (45)
where $\gamma_{\phi}\in\mathbb{R}_{>0}$ is the design parameter to be chosen properly. Furthermore, since $\text{sgn}(\phi)$ is a discontinuous function, an approximation of $\text{sgn}(\phi)\approx\tanh\Big{(}\dfrac{\phi}{\epsilon}\Big{)}$, where $\epsilon\in\mathbb{R}_{>0}$ is a negligible constant, can be evaluated to eliminate the discontinuity. As a consequence, the final form of the bank angle protection formulation for generating a roll rate command is presented in Eq. (46).
$u^{\star}=\operatorname*{argmin}_{p_{\text{GCAS}}\in\mathbb{R}}$ $\dfrac{1}{2}\big{(}p_{\text{GCAS}}-p_{\text{pilot}}\big{)}^{2}$ (46)
s.t. $-\tanh\Big{(}\dfrac{\phi}{\epsilon}\Big{)}[f(\phi)+g(\phi)p_{\text{GCAS}}]+\gamma_{\phi}b(\phi)\geq 0$
$p_{\text{min}}\leq p_{\text{GCAS}}\leq p_{\text{max}}.$
The constructed formulation enables the Auto-GCAS to generate a roll rate command ($p_{\text{GCAS}}$) that closely follows the roll rate command of pilot ($p_{\text{pilot}}$) while adhering to the constraint for ground collision avoidance, i.e. $\phi=0^{\circ}$. Additionally, the generated roll rate command, $p_{\text{GCAS}}$, must remain within the interval $p_{\text{min}}\hskip 4.26773ptp_{\text{max}}]$, considering the admissible and allowable roll rate limits depending on the aircraft. Since there does not exist an additional layer for the bank angle protection, the command $p_{\text{GCAS}}$ is equivalent to the command $p_{\text{cmd}}$.
Remark 4.1.
It is noteworthy that the bank angle must be within the interval, $\pi\leq\phi\leq\pi$, by definition.
@@
Figure 4: General framework of the proposed method: (1) Safety layer including GCAS and FEP, (2) CAS including NDI, (3) Control allocation, and (4) A/C dynamics.
The pilot commands $p_{\text{pilot}}$, $q_{\text{pilot}}$, and $r_{\text{pilot}}$, with roll and pitch rate commands subject to GCAS constraints. The pitch rate command is adjusted by the ECBF constraint, and $q_{\text{GCAS}}$ is generated if needed. This command is also limited by angle of attack and load factor protection, generating $q_{\text{cmd}}$ when required. Bank angle protection initiates a bank-to-level maneuver, generating $p_{\text{GCAS}}$ if $q_{\text{pilot}}-q_{\text{GCAS}}\neq 0$, indicating a recovery maneuver initiated; otherwise, the pilot’s roll rate command, $p_{\text{pilot}}$, is passed through.
5 Results
In order to calculate the adaptive ECBF gain vector, the optimization is performed under various conditions by meshing $\phi$, $\theta$, and $V_{T}$ within the intervals $150^{\circ}\hskip 4.26773pt150^{\circ}]$ with an increment of $\Delta\phi=50^{\circ}$, $60^{\circ}\hskip 4.26773pt-10^{\circ}]$ with an increment of $\Delta\theta=10^{\circ}$, and $200\hskip 4.26773pt350]$ with an increment of $\Delta V_{T}=30$m/s, respectively. The other flight states are randomly assigned to ensure the aircraft dives. Finally, the optimization is carried out at each design point, assuming the ground is flat and at zero level, i.e., $h_{\text{DTED}}=0$m while keeping $h_{\text{buff}}=100$ m, and the calculated gains $k_{1}$ and $k_{2}$ are scheduled as functions of $\phi$, $\theta$, and $V_{T}$. Moreover, the proposed Auto-GCAS design has been evaluated through three different assessments, (1) terrain collision avoidance scenarios, (2) an authority-sharing scenario, and (3) Monte Carlo simulations. Note that, for all assessments, the buffer height is set to 100 meters, and the center of gravity of the aircraft is set to $35\%\bar{c}$, which makes the aircraft statically unstable in the pitch axis.
5.1 Terrain Collision Avoidance Assessments
Two distinct terrain collision avoidance scenarios are conceptualized: one at low altitude with randomized state variables and another at high altitude with randomized states variables. The term “randomized state variables" encompasses aerodynamic angles, Euler angles, angular rates, and true velocity, with altitude being explicitly excluded. A default terrain model, synthesized using MATLAB®’s peaks function, is used to generate a smooth, synthetic terrain surface for visualization and conceptual validation of the barrier formulation; it does not represent real-world terrain. Furthermore, it is assumed that the pilot is incapacitated (e.g. experiencing blackout), resulting in the absence of control inputs, such that angular rate commands are maintained at $0^{\circ}/\text{s}$. Finally, a terrain scanning pattern should be designed to account for non-flat, mountainous terrain. However, since developing such a method is outside the scope of this study, a default rectangular scan pattern is employed, extending 750 meters ahead of the aircraft and 150 meters to each side (left and right).
5.1.1 Scenario-1: Low Altitude